Assume 75th percentile wage is 90$ per hour, while 50th percentile is 85$ per hour, and...

  1. Assume 75th percentile wage is 90$ per hour, while 50th percentile is 85$ per hour, and 25th percentile wage is 10$.
    1. Calculate w75/w50.
    2. Calculate w75/w25.
    3. Calculate w50/w25.
Answer #1

1. w75/w50 = 90/85 = 1.06

2. w75/w25 = 90/10 = 9

3. w50/w25 = 85/10 = 8.5.

Thanks!